# Hyperparameter Optimization for Randomized Algorithms: A Case Study for Random Features

### v2

@@ 3.3 Recasting Optimization as Inversion
To construct a suitable inverse problem in the form (35) so that the stochastic optimization of the RF empirical Bayes objective function (34) becomes amenable to EKI, we first take a fixed disjoint partition $\Omega\cup\Omega^{\textsf{c}}=\{1,\dots,N\}$, where $|\Omega|=\lceil rN\rceil$. In practice, we choose $r=0.8$. In what follows, the notation $X_{\Omega}$ denotes the set ${\mathopen{}\mathclose{{}\left\{x_{i}\in X\,\middle|\,i\in\Omega}\right\}}$ and similarly for $Y_{\Omega}$. The following items define the constituents of the proposed inversion framework:
@@
    namely, a contribution from the variability within $\mathcal{G}$ and from the noise perturbing the supervised training data. In (44), the covariance operation on $\mathcal{G}(u)$ is taken with respect to the random features and conditional on the realization of the noisy training data that underly the regression problem. The $2$ by $2$ identity matrix in the lower right block is needed to maintain consistency with the ideal empirical Bayes objective (34) in the large feature limit, as discussed next.
With these four ingredients in hand, we may now directly apply the EKI framework from Subsection 3.2. Iterating the EKI algorithm with small enough $\Delta t$ towards termination time $T=1$ will minimize the functional
${L}^{(\mathrm{EKI})}_{M}(u)\coloneqq\big{\lVert}\Gamma(u)^{-\frac{1}{2}}(z-\mathcal{G}(u))\big{\rVert}^{2}-2\log P(u),$ (45)
where $P(u)$ is the prior probability density of $u$. This is the weighted observable misfit added to a diminishing contribution from the prior on $u$ as the number of iterations increases. The prior contribution ensures that desired constraints such as positivity hold. The implied EKI objective function (45) is related to the ideal empirical Bayes cost (34) in the infinite feature limit. Indeed, in this limit, the term $\operatorname{Cov}_{\mu_{u}^{\otimes M}}(\mathcal{G}(u))$ tends to zero so that for large enough $M$, it holds that
@@
The cost ${L}^{(\mathrm{EKI})}_{M}$ (46) in this large feature limit is structurally similar to the cost ${L}^{(\mathrm{EB})}_{M}$ (34), but with two key differences. The data misfit in (46) is with respect to a validation partition $(X_{\Omega^{\textsf{c}}},Y_{\Omega^{\textsf{c}}})$ rather than on the full training data set $(X,Y)$. We observe in practice that this improves generalization accuracy. Since (46) arises from a Bayesian inverse problem, it also requires a prior to regularize the inversion, and hence optimization; this contribution appears in ${L}^{(\mathrm{EKI})}_{M}$ but not in ${L}^{(\mathrm{EB})}_{M}$. Although (46) is approximately valid for large numbers of random features, for efficiency purposes we solve the EKI training inverse problem with relatively few features in practice.
Remark 1 (Optimization as a Bayesian Inverse Problem)
We now detail additional practical considerations of the proposed stochastic inversion framework:
  • The law of $\mathcal{G}(u)$ and the prior distributions are assumed to be Gaussian. This assumption is not necessary, but it enables straightforward application of ensemble Kalman methods. The details of (not necessarily Gaussian) priors chosen for the numerical experiments appearing in the next section are provided in Appendix A.
  • In practice, $\Gamma(u)$ is approximated empirically by using an offline constant approximation $\Gamma(u)\approx\Gamma(u^{\dagger})$ at a fixed $u^{\dagger}$. In the following numerical results, $u^{\dagger}$ is taken as the mean of the prior. The effect of improved approximation of $\Gamma(u)$ is not considered in this investigation but remains an interesting direction for future work.
  • It is tempting to directly optimize the likelihood, i.e., define instead $\mathcal{G}(u)$ to be equal to $-\log P(z\,|\,u)$, and take $z=0$. This results in a one-dimensional optimization problem. Unfortunately, we report that this approach incurs large approximation errors in numerical experiments with input or output dimensions of even moderate size.
  • During the following experiments it was observed that under some problem realizations (typically around $10\%$ of cases), the norm penalty contribution $\frac{1}{M}\|\beta_{\Omega}^{(u)}\|^{2}$ in (46) would squash all posterior mean coefficients to zero. Therefore, this term was omitted from the objective function in practice.
  • Treating hyperparameter optimization for RFR as an black-box inverse problem also opens up the opportunity to replace the direct data misfit for the posterior mean with indirect observable-predictor pairs. Suitably balanced, this would allow for training of the embedded RFs within a complex system.
3.4 A Practical Kernel Structure
In this subsection, we instantiate our general framework with an implementable parametrization of the RF pair $(\varphi,\mu_{u})$. In all numerical experiments to follow, we employ real vector-valued random Fourier features (Rahimi and Recht, 2007) with Gaussian samples and an additional scaling parameter $\sigma>0$. Specifically, define for each $x\in\mathcal{X}\subset\mathbb{R}^{d}$ the $\mathbb{R}^{p}$-valued feature map
$\varphi(x;\theta)\coloneqq\sqrt{\sigma}\cos(\Xi x+B),\quad\text{where}\quad\theta=(\Xi,B)\sim\mathscr{D}$ (47)
@@ 3.4 A Practical Kernel Structure
The task is then to jointly learn the hyperparameters $(\sigma,U,S)$. We enforce that $S$ be diagonal, leading to a moderate number of total tunable parameters $r(dp+r)+1$. This is advantageous whenever $r\ll dp$.
There are many other options for parametrization beyond (vector-valued) random Fourier features of the form (47) and (48). For instance, one could parametrize a possibly sparse precision matrix instead of the Gaussian covariance $C$, or use other efficient representations such as FastFood kernels (Le et al., 2013; Yang et al., 2015). Experimentally, reducing the parameter count by imposing restrictive input-output separability or rank one structure of $C$ significantly limits expressivity and performance. For example, one pitfall would be to sample $\Xi\sim\mathsf{MatrixNormal}(0,C_{\mathrm{in}},C_{\mathrm{out}})$ and learn $C\simeq(C_{\mathrm{in}},C_{\mathrm{out}})\in\mathbb{R}^{d\times d}\times\mathbb{R}^{p\times p}$.
@@
Going beyond Gaussian samples, one can take inspiration from the scalar output $p=1$ case. Here, sampling $\Xi$ from a Student’s $t$ distribution leads to a limiting Matérn kernel; this follows from the fact that such densities are Fourier transform pairs. Similar results hold for Cauchy and Laplace distributions. In the $p>1$ case, one can still sample $\Xi$ from such non-Gaussian distributions on the reshaped $\mathbb{R}^{dp}$ space. The main hyperparameters that characterize these distributions are generalized lengthscale matrices $C\in\mathbb{R}^{dp\times dp}$, just as in the Gaussian case. Going beyond the Fourier setting, one approach is to work instead with random neural networks, where the feature map is a hidden neuron of the form $x\mapsto\varrho(\Xi x+B)$ for some activation function $\varrho(\cdot)$. These directions are left to future work.
@@ 4 Applications
To illustrate the role of the hyperparameter training, this section presents different applications that require the emulator to represent nonlocal features away from the training data. Global sensitivity analysis requires the posterior mean function to be accurate at quasi-random sampling of a $n$-dimensional cube (Subsection 4.1); to integrate a Lorenz 63 system, the posterior mean function must progress along, and attract towards, the Lorenz attractor (Subsection 4.2); for accelerated uncertainty quantification, the emulator mean and covariance predictions are used within a sampling method to explore the the support of a five-dimensional posterior distribution (Subsection 4.3). In all three applications, we observe moderate numbers of data that are perturbed by additive Gaussian noise with known covariance.
For comparison to the approach developed in the present investigation, two other emulators are presented, an untrained RF model using hyperparameters from the mean of the EKI initialization and a trained GP from a robust software package. The framework used to emulate smooth functions $\mathbb{R}^{p}\to\mathbb{R}$ with GPs is to use a radial basis function kernel with scaling parameter, nugget parameter, and learnable lengthscales in each of the $d$ input dimensions. To handle $p>1$ output dimensions, one uses an SVD decomposition in the output space to decorrelate, and then one trains $p$ decorrelated GP regressors in each dimension. Therefore a modest $d(p+2)$ parameters are found with gradient-based optimizers built into the chosen software packages. For our experiments we report results from SciKitLearn.jl v0.7. We also tried GaussianProcesses.jl v0.12.5 but found hyperparameter optimization to be too brittle to provide consistent comparisons.
@@ 4.1 Global Sensitivity Analysis
Global sensitivity analysis (GSA) seeks to attribute variance of model output to its input variables over a given domain (Saltelli, 2002). Practitioners are interested in either ranking variables by their direct contribution to the variance, computing $\mathsf{V}_{i}=\mathbb{V}(\mathbb{E}(f(x)|x_{i}))/\mathbb{V}(f(x))$, or residuals of the total variance when conditioning on all other variables, $\mathsf{TV}_{i}=(\mathbb{V}(f(x))-\mathbb{V}(\mathbb{E}(f(x)|x_{j\neq i})))/\mathbb{V}(f(x))$, where $x_{j\neq i}$ indicates all variables except $x_{i}$. Here $\mathbb{V}$ and $\mathbb{E}$ represent variance and expectation, respectively.
Two standard functions from the sensitivity analysis literature are used in the demonstration. These are chosen for their analytic representation of $\mathsf{V}_{i}$ and $\mathsf{TV}_{i}$. Furthermore, accurate empirical estimates of $\mathsf{V}_{i}$ and $\mathsf{TV}_{i}$ can be calculated on the cube using quasi-Monte Carlo methods. Functions are evaluated over points from a low discrepancy Sobol sequence (Sobol and Levitan, 1999) and are denoted as empirical estimates in our results.
@@ 4.1.1 Ishigami Function
$x\mapsto f(x;a,b)=(1+bx_{3}^{4})\sin(x_{1})+a\sin(x_{2})$ (49)
on $[-\pi,\pi]^{3}$ with $a=7$ and $b=0.1$ is considered. Data is taken as 300 evaluations of the Ishigami function from an 16000-long Sobol sequence, and Gaussian noise is added with variance $0.1^{2}$. AFter their hyperparameters are trained, the emulators are fit to the data and used to predict values over the entire Sobol sequence. This is how $\mathsf{V}_{i}$ and $\mathsf{TV}_{i}$ are calculated empirically.
Figure 1 shows the results from a single trial. Row 1 shows slices of the true Ishigami function against each variable evaluated on the Sobol sequence. Rows 2-4 show the noisy data (red) used for learning and the emulator predictions over the Sobol sequence (blue). Row 2 corresponds to the untrained RF, Row 3 is the prediction from a trained GP, and Row 4 shows results of the EKI-trained RF(500 features). The successful training runs show good fit to the true function.
Table 2 displays the calculated Sobol indices of the trained emulators; accuracy here is a good metric for GSA. To gauge the robustness of randomized training algorithms, the training was repeated over 20 different initial samples, and Table 2 displays the mean and standard deviation of these trials. For both $\mathsf{V}_{i}$ and $\mathsf{TV}_{i}$, both GP and RF produce close values, with RF having variations at most 5% of the true sensitivity, and true values always lie within 2 standard deviations of the RF emulator variability.
Figure 1: Learning the $\mathbb{R}^{3}\to\mathbb{R}$ Ishigami function from 300 noisy samples. In orange, Sobol samples of the true function; in red, 300 noisy observations; in blue, Sobol samples of the emulated functions. Row 2 is sampled from an RF emulator without hyperparameter training, row 3 are samples from a trained GP emulator, and row 4 is the RF approximation (full rank) using our hyperparameter optimizer.
Table 2: First order global sensitivity analysis of the Ishigami function. Displayed are the analytic values, the empirical approximation from evaluating the true function at 16,000 Sobol indices, and the empirical approximations obtained from evaluating the trained GP and RF emulators at the Sobol indices. The final column estimates a mean and standard deviation of the emulators from repeating the training procedure 20 times on the same data.
Sobol Index Analytic Empirical Trained GP Trained RF
@@
Further details of the implementation of this experiment relating to EKI configuration and convergence are provided in Appendix A.3. The algorithm is robust to the data seeding, the number of features used for optimization and prediction, and to various modifiers of the optimization algorithm.

### v4

@@ 3.3 Recasting Optimization as Inversion
We construct a suitable inverse problem in the form (35) so that the stochastic optimization of the RF empirical Bayes objective function (34) becomes amenable to EKI. Take $\Omega\subseteq[N]$ and $\Omega^{\textsf{c}}\subseteq[N]$ to be training and validation index sets, respectively. In what follows, the notation $X_{\Omega}$ denotes the set ${\mathopen{}\mathclose{{}\left\{x_{i}\in X\,\middle|\,i\in\Omega}\right\}}$ and similarly for $Y_{\Omega}$. The following items define the constituents of the proposed inversion framework:
@@
    In particular, the square root in the third component of (41) is well-defined.
@@
    namely, a contribution from the variability within $\mathcal{G}$ and from the noise perturbing the labeled training data. In (44), the covariance operation on $\mathcal{G}(u)$ is taken with respect to the random features and is conditional on the realization of the noisy training data that underlie the regression problem. The $2$ by $2$ identity matrix in the lower right block is needed to maintain consistency with the ideal empirical Bayes objective (34) in the large feature limit, as discussed next.
With these four ingredients in hand, we may now directly apply the EKI framework from Subsection 3.2. Iterating the EKI algorithm with small enough $\Delta t$ towards termination time $T=1$ will minimize the functional ${L}^{(\mathrm{EKI})}_{M}$ defined by
${L}^{(\mathrm{EKI})}_{M}(u)\coloneqq\big{\lVert}\Gamma(u)^{-\frac{1}{2}}\bigl{(}z-\mathcal{G}(u)\bigr{)}\big{\rVert}^{2}-2\log P(u),$ (45)
where $P(u)$ is the prior probability density of $u$; cf. (38). This is the weighted observable misfit added to a diminishing contribution from the prior on $u$ as the number of iterations increases. The prior contribution ensures that desired constraints such as positivity hold. The implied EKI objective function (45) is related to the ideal empirical Bayes cost (34) in the infinite feature limit. Indeed, in this limit, the term $\operatorname{Cov}_{\mu_{u}^{\otimes M}}(\mathcal{G}(u))$ tends to zero so that for large enough $M$, it holds that
@@
The terms on the right hand side of (46) in this large feature limit are structurally similar to those in the cost ${L}^{(\mathrm{EB})}_{M}$ (34) when $\Omega=\Omega^{\textsf{c}}=[N]$. In practice, we prefer instead a construction similar to cross-validation: take a fixed disjoint partition $\bigcup_{k=1}^{\mathsf{K}}\Omega_{k}=[N]$, where $|\Omega_{k}|=N/\mathsf{K}$ and $\mathsf{K}$ divides $N$. We define a validation partition for an index $j\in[\mathsf{K}]$ as $\Omega^{\textsf{c}}\coloneqq\Omega_{j}$; the remaining indices form the training partition $\Omega\coloneqq\bigcup_{k\neq j}\Omega_{k}$. Repeating this procedure by cycling through $j\in[\mathsf{K}]$, we may incorporate up to $\mathsf{K}$ partitions by stacking for each $j$ the corresponding data $z$ for each validation partition and comparing this to stacked forward map evaluations $\mathcal{G}(u)$ built from the associated training partition, plus additive Gaussian noise with block diagonal covariance matrix having $\Gamma$ as each block. The introduction of these validation partitions improves computational efficiency, which scales with $|\Omega^{\textsf{c}}|\leq N$ instead of $N$. With even two such partitions, we also observe improved convergence of the optimizer and improved accuracy in experiments. We hypothesize that cross-validation can improve the landscape of the objective function but leave further investigation to future work. Since (46) arises from a Bayesian inverse problem, it also requires a prior to regularize the inversion, and hence optimization; this contribution appears in ${L}^{(\mathrm{EKI})}_{M}$ and in (46) but not in ${L}^{(\mathrm{EB})}_{M}$. Although (46) is approximately valid for large numbers of random features, for efficiency purposes we minimize it with relatively few features in practice. [^8] The effect of varying the number of features during optimization is explored in Appendix A.3.1.
3.4 A Practical Feature Distribution Structure
In this subsection, we instantiate our general framework with an implementable parametrization of the RF pair $(\varphi,\mu_{u})$. In all numerical experiments to follow, we employ real vector-valued random Fourier features (Rahimi and Recht, 2007) with Gaussian samples and an additional scaling parameter $\varsigma>0$. Specifically, define for each $x\in\mathcal{X}\subset\mathbb{R}^{d}$ the $\mathbb{R}^{p}$-valued feature map
$\varphi(x;\theta)\coloneqq\sqrt{\varsigma}\cos(\Xi x+B),\quad\text{where}\quad\theta=(\Xi,B)\sim\mathscr{D}$ (47)
@@ 3.4 A Practical Feature Distribution Structure
The task is then to jointly learn the hyperparameters $(\varsigma,U,S)$. We enforce that $S$ be diagonal, leading to a moderate number of total tunable parameters $r(dp+r)+1$. This is advantageous whenever $r\ll dp$. We refer to the preceding construction as a nonseparable feature distribution with rank $r$. We also consider $\Xi\sim\mathsf{MatrixNormal}(0,C_{\mathrm{in}},C_{\mathrm{out}})$ (Dutilleul, 1999) and tune the pair $(C_{\mathrm{in}},C_{\mathrm{out}})\in\mathbb{R}^{d\times d}\times\mathbb{R}^{p\times p}$. This pair of covariances can be identified with a larger covariance $C\in\mathbb{R}^{dp\times dp}$ with Kronecker structure. This construction is referred to as a separable feature distribution with rank $(r_{\mathrm{in}},r_{\mathrm{out}})$ and leads to $r_{\mathrm{in}}(d+r_{\mathrm{in}})+r_{\mathrm{out}}(p+r_{\mathrm{out}})+1$ total tuneable parameters. A suitable rank is selected to produce emulators with minimal $L^{2}$ errors on a held-out test set in experiments.
@@
There are many other options for parametrization beyond (vector-valued) random Fourier features of the form (47) and (48). For instance, one could parametrize a possibly sparse precision matrix instead of the Gaussian covariance $C$, or use other efficient representations such as FastFood kernels (Le et al., 2013; Yang et al., 2015). Going beyond Gaussian samples, one can take inspiration from the scalar output $p=1$ case. Here, sampling $\Xi$ from a Student’s $t$ distribution leads to a limiting Matérn kernel; this follows from the fact that such densities are Fourier transform pairs. Similar results hold for Cauchy and Laplace distributions. In the $p>1$ case, one can still sample $\Xi$ from such non-Gaussian distributions on the reshaped $\mathbb{R}^{dp}$ space. The main hyperparameters that characterize these distributions are generalized lengthscale matrices $C\in\mathbb{R}^{dp\times dp}$, just as in the Gaussian case. Going beyond the Fourier setting, one approach is to work instead with random neural networks, where the feature map is a hidden neuron of the form $x\mapsto\varrho(\Xi x+B)$ for some activation function $\varrho(\cdot)$. These directions are left to future work.
3.5 Current Limitations of the Methodology
We now detail additional practical considerations of the proposed stochastic inversion framework as well as some of its limitations. First, we highlight that the law of $\mathcal{G}(u)$ and the prior distributions are assumed to be Gaussian in our formulation. This assumption is not strictly necessary, but it enables straightforward application of ensemble Kalman methods.
Though RFR has admirable scaling, the objective function (34) scales naïvely as $(Np)^{2}$. This significantly slows down workflows involving high-dimensional spaces or large data sets. Therefore, controlling the computational cost necessitates further approximations at implementation time. One example, already discussed in Section 3.3, is to evaluate the objective on a validation partition of cardinality less than or equal to the total sample size $N$. Another approximation is to avoid computing $u\mapsto\Gamma(u)$ in (44) at all $u$ (requiring the covariance of (41) under the random features) by using an offline constant approximation $\Gamma(u)\approx\Gamma(u^{\dagger})$ at a fixed $u^{\dagger}$. In the following numerical results, $u^{\dagger}$ is taken as the mean of the hyperparameter prior. The effect of improved approximation of $\Gamma(u)$ is not considered in the current paper but remains an interesting direction for future work.
Finding local optima naturally has dependence on the initial condition. For EKI, this arises from the prior used to create the initial ensemble. There is much debate on prior selection (Gelman et al., 2017), particularly in high-dimensional and nonphysical systems. The heuristic adopted in the present work is to choose weakly informative priors over the hyperparameters. We use the same prior structure across all of the forthcoming numerical experiments. Appendix A.1 contains the details. Moreover, it is tempting to directly optimize the likelihood, i.e., define instead $\mathcal{G}(u)$ to be equal to the scalar value $-\log P(z\,|\,u)$ and take $z=0$ as the “observable”. This results in a one-dimensional optimization problem. Unfortunately, we report that this approach incurs large approximation errors in numerical experiments with input or output dimensions of even moderate size. The more costly vector-valued formulation in Subsection 3.3 is preferred.
@@ 4 Applications
To illustrate the role of the hyperparameter tuning, this section presents different applications that require the machine learning emulator to represent nonlocal features away from the training data. Global sensitivity analysis requires the posterior mean function to be accurate at quasi-random samples from the $d$-dimensional cube (Subsection 4.1); to integrate a Lorenz 63 system, the posterior mean function must progress along, and attract towards, the Lorenz attractor (Subsection 4.2); for accelerated uncertainty quantification, the emulator mean and covariance predictions are used within a sampling method to explore the support of a five-dimensional posterior distribution (Subsection 4.3). In all three applications, the data consist of a moderate number of observations that are perturbed by additive Gaussian noise with known covariance matrix.
For comparison to the approach developed in the present paper, two other emulators are presented, an untuned RF model using hyperparameters from the mean of the EKI initialization and a tuned GP from a robust software package. To emulate smooth functions $\mathbb{R}^{d}\to\mathbb{R}$ (i.e., $p=1$) with GPs, we use a radial basis function kernel with scaling parameter, nugget parameter, and learnable lengthscales in each of the $d$ input dimensions. To handle $p>1$ output dimensions, we use a singular value decomposition (SVD) in the output space to decorrelate the coordinates, and then fit and tune $p$ decorrelated GP regressors in each dimension. This corresponds to a single vector-valued GP regressor with a diagonal matrix-valued kernel. Therefore, a modest $d(p+2)$ parameters are found with gradient-based optimizers built into the chosen software packages. In our experiments we report results from SciKitLearn.jl v0.7. We also tried GaussianProcesses.jl v0.12.5, but found hyperparameter optimization to be too brittle to provide consistent comparisons.
Design Choices.
To support many of the heuristics in the upcoming experiments, we provide supplementary investigations in Appendix A for each application. These include details about the EKI configuration, objective function configuration, and convergence plots over different random initializations. Some other specific studies we include are run times and accuracy with respect to input dimension size (Appendix A.3) and numbers of features (Appendix A.3.1), and accuracy over different feature distribution ranks and covariance structures in the multi-output setting (Appendix A.4.1 and A.5).
@@ 4.1 Global Sensitivity Analysis
Global sensitivity analysis (GSA) seeks to attribute the variance of model output to its input variables over a given domain (Saltelli, 2002). Practitioners are interested in either ranking variables by their direct contribution to the variance, computing $\mathsf{V}_{i}=\mathbb{V}(\mathbb{E}(f(x)\,|\,x_{i}))/\mathbb{V}(f(x))$, or residuals of the total variance when conditioning on all other variables, $\mathsf{TV}_{i}=(\mathbb{V}(f(x))-\mathbb{V}(\mathbb{E}(f(x)\,|\,x_{j\neq i})))/\mathbb{V}(f(x))$. Here $x_{j\neq i}$ indicates all variables except $x_{i}$ and $\mathbb{V}$ represents the variance operator.
Two standard functions from the sensitivity analysis literature are used in the demonstration. These are chosen for their analytic representation of $\mathsf{V}_{i}$ and $\mathsf{TV}_{i}$. Furthermore, accurate empirical estimates of $\mathsf{V}_{i}$ and $\mathsf{TV}_{i}$ can be calculated on the cube using quasi-Monte Carlo methods. Functions are evaluated over points from a low discrepancy Sobol sequence (Sobol and Levitan, 1999) and are referred to as empirical estimates in our results.
@@ 4.1.1 Ishigami Function
$x\mapsto f(x;a,b)\coloneqq(1+bx_{3}^{4})\sin(x_{1})+a\sin(x_{2})$ (49)
on $[-\pi,\pi]^{3}$ with $a=7$ and $b=0.1$ is considered. Data is taken as 300 evaluations of the Ishigami function from a length 16000 Sobol sequence, and Gaussian noise is added with variance $0.1^{2}$. After their hyperparameters are tuned, the emulators are fit to the data and used to predict values over the entire Sobol sequence. This is how $\mathsf{V}_{i}$ and $\mathsf{TV}_{i}$ are calculated empirically.
Figure 1: Learning the $\mathbb{R}^{3}\to\mathbb{R}$ Ishigami function from $300$ noisy samples. In orange, Sobol samples of the true function; in red, $300$ noisy observations; in blue, Sobol samples of the emulated functions. Row 2 is sampled from an RF emulator without hyperparameter tuning, Row 3 is sampled from a tuned GP emulator, and Row 4 is sampled from the RF approximation (using a nonseparable feature distribution with full rank covariance) using our hyperparameter optimizer.
Figure 1 shows the results from a single trial. Row 1 shows slices of the true Ishigami function against each variable evaluated on the Sobol sequence. Rows 2–4 show the noisy data (red) used for learning and the emulator predictions over the Sobol sequence (blue). Row 2 corresponds to the untuned RF, Row 3 is the prediction from a tuned GP, and Row 4 shows results of the EKI-tuned RF with $500$ features. The successful trials show that the tuned models fit the true Ishigami function well.
Table 2: First order global sensitivity analysis of the Ishigami function. Displayed are the analytic values, the empirical approximation from evaluating the true function at $16000$ Sobol indices, and the empirical approximations obtained from evaluating the tuned GP and RF emulators at the Sobol indices. The final column estimates a mean and standard deviation (in parentheses) of the emulators from repeating the stochastic tuning procedure $20$ times on the same data.
Sobol Index Analytic Empirical Tuned GP Tuned RF
@@
To quantify the accuracy of GSA, Table 2 displays the calculated Sobol indices of the tuned emulators. The tuning was repeated over $20$ different initial samples to gauge the robustness of randomized hyperparameter learning algorithms. Table 2 shows the mean and standard deviation of these trials. For both $\mathsf{V}_{i}$ and $\mathsf{TV}_{i}$, both GP and RF produce close values. In particular, RF has variations of at most 5% of the true sensitivity and the true values always lie within two standard deviations of the RF emulator variability.
Further details about the implementation of this experiment relating to EKI configuration and convergence are provided in Appendix A.3. The algorithm is robust to the data seeding, the number of features used for optimization and prediction, and to various modifiers of the optimization algorithm.